OpenStudy (anonymous):

find the values for a and b such that lim x->0 (((sqrt a+bx)-5)/x)=4

OpenStudy (anonymous):

Is this right?\[ \lim_{x\to 0}\frac{\sqrt{a+bx}-5}{x}=4 \]

OpenStudy (anonymous):

Try multiplying by the conjugate.

OpenStudy (anonymous):

I've done that and got: a+bx-25/x((sqrt a+bx)+5x)=4

OpenStudy (anonymous):

You get\[ \lim_{x\to 0}\frac{a+bx-25}{x(\sqrt{a+bx}+5)} \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm not sure of the next step

OpenStudy (anonymous):

You want the \(x\)s to cancel out. So you want \(a-25=0\)

OpenStudy (anonymous):

How would you get the bx's to cancel in order to get a-25

OpenStudy (anonymous):

Okay so you want\[ a+bx-25 = bx \]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

That way the numerator is divisible by \(x\).

OpenStudy (anonymous):

What did you do to get to that step from a+bx-25/(x(sqrt.a+bx)+5)=4

OpenStudy (anonymous):

What cancels to get a+bx-25=bx

OpenStudy (anonymous):

Okay so look at the denominator: \[ x(\sqrt{a+bx}+5) \]We doon't want it ot be \(0\).

OpenStudy (anonymous):

Exactly, what can be done about the denominator

OpenStudy (anonymous):

So if the \(x\) cancel, we have: \[ \lim_{x\to0}\frac{b}{\sqrt{a+bx}+5} \]

OpenStudy (anonymous):

Wouldn't that still yield 0 because an x is on the bottom?

OpenStudy (anonymous):

As \(x\to0\) we have \[ \frac{b}{\sqrt{a+b(0)}+5} = \frac{b}{\sqrt{a}+5} \]

OpenStudy (anonymous):

Where did the a go on the numerator as the x's cancelled

OpenStudy (anonymous):

Okay, I need to breath in and breath out and summon my patience here.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio Okay so you want\[ a+bx-25 = bx \] \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @wio You want the \(x\)s to cancel out. So you want \(a-25=0\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

When you get b/((sqrt a)+5)=4 you'll have to solve for "b" by multiplying by ((sqrt a)+5)) on both sides to get b=((sqrt a)+5)+4. After simplification how do you solve for b as b=(sqrt a)+9?

OpenStudy (anonymous):

We have two equations: \[ a-25=0\\ b=4(\sqrt{a} + 5) \]

OpenStudy (anonymous):

Sorry it's not +4 it's x4

OpenStudy (anonymous):

So a=25 and b=40?

OpenStudy (anonymous):

Yes. Now test it.

OpenStudy (anonymous):

Thank you! I understand the process, but how did you know to look for a+bx-25=bx?

OpenStudy (anonymous):

Because you want an \(x\) in the numerator and an \(x\) in the denominator to cancel.

OpenStudy (anonymous):

So the numerator has to be divisible by \(x\).

OpenStudy (anonymous):

So it has to be \(bx\)

OpenStudy (anonymous):

Okay, thank you wio!